### v1

# Inhibition of bacterial growth by antibiotics

Abstract
Growth in bacterial populations generally depends on the environment (availability and quality of nutrients, presence of a toxic inhibitor, product inhibition..). Here, we build a general model to describe the action of a bacteriostatic antibiotic, assuming that this drug inhibits essential autocatalytic cycles involved in the cell metabolism. The model can describe various types of antibiotics and confirms the existence of two distinct regimes of growth-dependent susceptibility, previously identified only for ribosome targeting antibiotics. Interestingly, below a certain threshold, a coexistence of two values of the growth rate is possible, which has also been observed experimentally.
autocatalytic networks, environment, antibiotics, growth
Introduction
The emergence of antibiotic resistance, which often occurs under changing levels of antibiotics is a major concern for human health [1]. In an important class of antibiotics, called bacteriostatic antibiotics [2], the drug does not induce death directly, but only renders some essential process in the cell metabolism less efficient or inactive [3, 4, 5, 6, 7, 8]. For these antibiotics, it thus appears essential to properly model cell metabolism and cell growth in order to better understand the action of antibiotics [9, 10, 11, 12].
In the field of bacterial growth, the experimental discovery of growth laws in the last decade [13, 12, 14, 15] represents a major step forward in our understanding of cell growth. These growth laws result from mass conservation and flux balance at steady-state. The first growth law has been derived using a comprehensive model of the cell metabolism based on the coupling of essential autocatalytic cycles, such as the cycle of ribosome production and that of RNA polymerase production [16]. This approach has also been used recently to formulate predictions about the interplay between cellular growth rate and mRNA abundances [17].
While predictions about the action of RNA-polymerase targeting antibiotics have also been derived from this framework, the full consequences for the inhibition of growth by a general antibiotics have not. In particular, Ref. [16] does not discuss the second growth law, nor the two modes of action of antibiotics, called reversible and irreversible binding regimes of antibiotics. This distinction is quite important in practice because for reversible binding, faster growth in the absence of the drug leads to an increased susceptibility, while the opposite is true for irreversible binding [12]. Further, the coexistence of two values of growth rate (growth rate bistability [18]) may occur below a certain threshold. At the moment, it is not known whether these behaviors should be expected for all types of antibiotics.
To summarize, we believe that the inhibition of bacterial growth by antibiotics has not been considered from a sufficiently general point of view, which is the approach we develop in the present paper. By building on Ref. [16], we develop a framework to describe the inhibition of bacterial growth by bacteriostatic antibiotics based on a model of cell metabolism. We first present our model, explore some of its consequences and extensions, and then show that it can describe successfully the dependence of the growth rate as function of the concentration of antibiotics for a wide range of different antibiotics. Further, we show that in a some limit, our general autocatalytic model allows to recover the equations of [12].
Model
We model the cell metabolism as two coupled autocatalytic cycles, in which one cycle describes the production of ribosomes, while the other describes RNA-polymerase production. These two autocatalytic cycles are coupled because ribosomes are necessary to synthesize RNA-polymerase protein subunits and vice-versa for ribosomes. To that basic model, we then add interaction with bacteriostatic antibiotics, as shown in the chemical network of Fig.1a: $B_{1}$ represents the number of active ribosomes; $C_{1}$ the number of active RNA polymerases; similarly $B_{2},...,B_{N-1}$ and $C_{2},...,C_{K-1}$ are the abundances of intermediates involved in the assembly of ribosomes and RNA polymerases respectively, $B_{N}$; $C_{K}$ are the abundances of fully assembled but resting ribosomes/RNA polymerases respectively, $R_{N},R_{K}$ are the abundances of building blocks needed to build $B_{N}$ and $C_{K}$.
Figure 1: (a) Scheme of coupled autocatalytic networks interacting with a toxic agent. The straight line linking two arrows represents a coupling through a min function [19, 20]. (b) Illustration of the growth laws when varying either the amount of antibiotics or the nutrient quality linked to pre-exposure growth rate $\lambda_{0}$ displayed on the right scale.
We suppose that "toxic" inhibiting agents in numbers $A$ can bind to one of the autocatalysts (chosen here to be $B_{1}$ for simplicity) with a rate $k_{on}$ and unbind with a rate $k_{off}$, proportionally to the relative abundance of antibiotics in the cell [12, 21]. We denote $B_{1,u}$ the abundance of unbound ribosomes and $B_{1,b}$ the abundance of bound ribosomes. The binding only occurs inside the cell, viewed as a compartment, in which antibiotics enter with rate $P_{in}$ and exit with rate $P_{out}$ (thanks to diffusion by passive transport or through pores by active transport) [22, 23]). The concentration of antibiotics outside the cell is $a_{ex}$. We use fractions measured with respect to the total number of mature individuals $B_{tot}=B_{1,u}+B_{1,b}+B_{N}$, assuming the total density of ribosomes remains constant in the cell of volume [24].
We rely on Leontief’s approach [19], or Liebig’s model in ecology [20], in which the rates of reactions involving two complementary resources are set by the limiting quantity among the two using a minimum function [16]. We denote $\tau_{life}$ the life time of mature individuals $B_{N},B_{1,u},B_{1,b}$, and we assume that we can neglect the inverse lifetimes of the intermediates $B_{2},...,B_{N-1}$. The names of the rates are self-explanatory and correspond to the transitions displayed in Fig.1a. In the following, we assume the cycle targeted by the toxic agent becomes limiting. Consequently, we can isolate this cycle and study its growth, because it restricts the growth of the rest of the network; the influence of the inhibition of the first cycle on the second cycle is studied in the Supplementary Material [21].
Due to balance growth of the cell, all species grow at the same rate $\lambda=d\ln{{\mathcal{N}}}/dt$, where $\mathcal{N}$ is typically the number of ribosomes or RNA-polymerases… One can then combine the equations of the model to obtain a linear matrix equation for the sub-populations of ribosomes only, without explicit dependence on antibiotics, and a self consistent equation for the growth rate $\lambda$ of the whole cycle (see Supplementary Material [21]). In the absence of inhibitors, the pre-exposure or basal growth rate is $\lambda_{0}$, which corresponds to the normal behaviour of the cell. As the concentration of antibiotics increases, the growth rate always decreases below this basal growth rate.
A key quantity is the fraction $Q(\lambda)=B_{1,u}/B_{tot}$, which takes the form of a polynomial :
$Q(\lambda)=\frac{1}{k_{B,1}}\left(1+\frac{\lambda}{k_{B,2}}\right)\times\ldots\times\left(1+\frac{\lambda}{k_{B,N-1}}\right)\left(\lambda+\frac{1}{\tau_{life}}\right).$ (1)
This polynomial simplifies in the limit of "fast assembly", which corresponds to $k_{B,2},...,k_{B,N-1}\gg\lambda$. In this case, we find the linear behaviour
$\frac{B_{1,u}}{B_{tot}}\simeq\frac{1}{k_{B1}}\left(\lambda+\frac{1}{\tau_{life}}\right).$ (2)
This linear increase between the fraction of unbound ribosomes and $\lambda$ is the first growth law [15, 13, 12, 16], which has been recovered from an autocatalytic cycles description of the cell in [16]. The law implies that by increasing the nutrient quality, which amounts to increase $\lambda_{0}$ by acting on assembly rates at external antibiotics concentration $a_{ex}=0$ with all other parameters fixed, we increase the growth rate and the fraction of mature ribosomes as illustrated in Fig. 1b.
The second growth law [13] predicts an increase of the growth rate together with a decrease in the activated ribosome fraction $(B_{1,u}+B_{1,b})/B_{tot}$ when the level of inhibition is increased. With our formalism, we obtain (see Supplementary Material [21]), in the limit of fast assembly, fast activation, long lifetime with a single intermediate step ($N=3$) :
$\frac{B_{1,u}+B_{1,b}}{B_{tot}}=1-\frac{\lambda}{k_{B,3}}.$ (3)
As shown on Fig.1b, by increasing the external concentration of antibiotics $a_{ex}$ with all other parameters fixed, we indeed predict an increase in the fraction of mature ribosomes and a decrease of the growth rate.
It is important to appreciate that the first and the second growth laws are derived from the model, they are not introduced as phenomenological constraints as done in Ref. [12]. It is also interesting to note that the solid blue curve in Fig. 1, which represents the limiting value of the ribosome fraction when $a_{ex}=0$ is not a linear function of $\lambda$, but a non-linear function with a negative curvature. This curvature was not expected in the original work on the growth laws [15] but it is also present in more detailed models of cell metabolism that include stochastic effects [25, 26].
We now explore further consequences of our formalism. For ribosomes, we can expect a long lifetime $\lambda\gg\frac{1}{\tau_{life}}$, a small resting rate $\lambda\gg k_{B,4}$, fast assembly and fast activation $\lambda\ll k_{B,3}$ [16]. These conditions translate to $\frac{1}{\tau_{life}},k_{B4}\ll\lambda_{0},k_{B1}\ll k_{B2},...,k_{B,N}$, yielding $\lambda_{0}\simeq k_{B1}$. In this limit, we can simplify our self-consistent equation for the growth rate :
$\frac{P_{in}a_{ex}}{\left(\frac{k_{B1}}{k_{on}}\frac{\lambda+P_{out}}{\lambda}+\frac{\lambda}{\lambda+k_{off}}\right)}\simeq\left(1-\frac{\lambda}{\lambda_{0}}\right)\left(\lambda+k_{off}\right).$ (4)
This equation is similar to that found in [12], which sets the growth rate of a bacteria in the presence of a bacteriostatic antibiotic. With the additional assumption of fast binding $\lambda_{0}\ll k_{on}$, the possible values of the growth rate are roots of a polynomial, from which it is possible to recover the reversible and irreversible limits of antibiotics binding [12]. In particular we find (see Supplementary Material [21]) a condition defining the transition from irreversible to reversible.
Reversible limit
The reversible limit corresponds to a strong outflux of toxic agents and unbinding rate. We find an equation setting the growth rate in this limit (see Supplementary Material [21]):
$Q(\lambda)=\frac{1}{1+\frac{K_{D}P_{in}}{P_{out}}a_{ex}}.$ (5)
In the limit $P_{out},k_{off}\gg\lambda_{0}$, the expression of the growth rate is that of Ref. [12], $\lambda=\frac{\lambda_{0}}{1+\frac{\lambda_{0}P_{in}a_{ex}}{k_{B,1}K_{D}P_{out}}}$, where $K_{D}$ is the dissociation constant $k_{off}/k_{on}$.
Irreversible limit
On the contrary, the irreversible limit corresponds to negligible outflux and unbinding rate compared to the influx of toxic agents and binding rate. Then, we obtain a different equation setting the growth rate (see Supplementary Material [21]):
$Q(\lambda)=1+\frac{P_{in}a_{ex}}{\lambda}.$ (6)
This equation typically has several solutions depending on the order of the polynomial $Q(\lambda)$. In the limit $P_{out},k_{off}\ll k_{on},\lambda_{0}$, we recover $\lambda=\frac{\lambda_{0}}{2}\left(1+\sqrt{1-\frac{4P_{in}a_{ex}}{\lambda_{0}}}\right)$ [12].
Interestingly, the self-consistent equation for the growth rate obtained within the autocatalytic framework (see Supplementary Material [21]) has two solutions in the irreversible limit with fast assembly, leading to two separate branches of solutions for $\lambda$. A first solution remains close to 00, corresponding to a non-growing cell. A second one is larger but exists only until a given concentration of inhibitors is reached, above which the system jumps on the other branch, and the growth rate vanishes as shown in Fig. 3a. In experiments, in the irreversible case, the system usually starts from $\lambda_{0}$ and the growth rate decreases as the concentration of inhibitors increases, until the discontinuity where the growth rate jumps on the second branch and vanishes. This growth bistability happens above a threshold, which can be determined by an implicit equation [21]. Such a phenomenon has been predicted in other theoretical work [12, 27], and it has also been observed experimentally [18, 26].
We have tested our model on a number of bacteriostatic antibiotics [2, 3, 6]: Chloramphenicol inhibits ribosome production by binding to ribosomes, preventing them from transcribing new proteins; Rifampicin targets RNA-polymerase by binding to RNA-polymerase [28, 29]; Kanamycin, Streptomycin, Chloramphenicol and Erythromycin target the ribosomal autocatalytic cycle [30, 3, 5, 7]; and finally Triclosan targets the synthesis of fatty acids [31, 32, 33], thus affecting the building of bacterial membranes [16].
Figure 2: Comparison with experiments for two bacteriostatic drugs, namely (a) Chloramphenicol (data from [8]) and (b) Kanamycin (data from [12]). The solid line shows the growth rate as a function of the fraction of inhibitors, while the dotted line shows a measure of the risk faced by the cell defined in the text. The data were fitted by constraining the parameters as explained is Supplementary Material [21].
In Fig.2, we show the normalized growth rate $\lambda/\lambda_{0}$ as function of the concentration of antibiotics only for Chloramphenicol and Kanamycin, the plots for the other antibiotics are shown in Supplementary Material [21].
In [16], the effects of Triclosan and Rifampicin were explained by adding Hill functions heuristically to describe saturation effects in the cycle. In contrast here, we provide an explicit expression for the dependence of the growth rate on the fraction of bacteriostatic antibiotics without such an assumption. The fact that we are able to describe a large panel of bacteriostatic antibiotics suggests that these antibiotics can indeed be depicted as inhibitors affecting essential cellular autocatalytic cycles despite their different mechanisms. Note that we recover different concavities in Fig.2, which correspond to the two distinct regimes of cellular response to the antibiotics previously identified for ribosome-targeting antibiotics [12]: the reversible limit where the outflux of antibiotics compensates the influx of the latter, and the irreversible limit where antibiotics bind quickly to autocatalysts, resulting in an accumulation of bound, inhibited individuals.
All these antibiotics are bacteriostatic agents, which slow growth but do not to induce death directly [11]. However, if the inhibition is too strong, processes that are necessary for survival cannot be satisfied and cell death can be induced in this way [34, 9]. To quantify this, we have introduced a measure of the risk faced by the cell, which we define as the fraction of bound active individuals $B_{1,b}$ with respect to unbound active individuals $B_{1,u}$ (see [21] for more details). The main interest of this notion is that it is independent of the type of action of the antibiotic and can be used to compare the efficiency of different antibiotics. This risk shown in Fig.2 as dashed lines is an increasing function of the concentration of antibiotics.
Half-inhibition concentration
The half-inhibition concentration $IC_{50}$ is defined as the concentration of toxic agent at which the growth rate is half its initial value. This is a measure of the sensitivity of the system to external stress, the higher it is, the more resistant is the system to inhibitors. We obtain an explicit expression for $IC_{50}$ in the limit of long lifetime and fast assembly, when the network contains an arbitrary number of steps $N$ (see [21] for details). If we can lump all intermediates into just one ($N=3$), we obtain
$\frac{IC_{50}}{IC_{50}^{*}}=\frac{1}{2}\left(\left(\frac{\lambda_{0}^{*}}{\lambda_{0}}+2K_{D}\frac{\lambda_{0}}{\lambda_{0}^{*}}\right)\left(1+\frac{\lambda_{0}}{2k_{off}}\right)+\frac{\lambda_{0}}{\lambda_{0}^{*}}\right),$ (7)
where we have rescaled the half-inhibition concentration by a typical concentration $IC_{50}^{*}$ and the basal growth rate by a typical value $\lambda_{0}^{*}$ [21]. Note that this expression does not depend only on the ratio $\lambda_{0}/\lambda_{0}^{*}$ but also on $\lambda_{0}$ (itself defined by the parameters of the system). The rescaled half-inhibition concentration as a function of the rescaled basal growth rate in this limit is the convex function shown in Fig.3b. Remarkably, this function allows to collapse the measurements of many types of antibiotics. We reproduce in this figure experimental data from Ref. [12].
Additionally, we find in the limit of long lifetime, fast binding, fast assembly, and with $k_{off}\gg\lambda_{0}$, the rescaled half-inhibition concentration is essentially $\frac{IC_{50}}{IC_{50}^{*}}\simeq\frac{1}{2}\left(\frac{\lambda_{0}^{*}}{\lambda_{0}}+\frac{\lambda_{0}}{\lambda_{0}^{*}}\right)$.
Figure 3: (a) Normalized growth rate versus the normalized antibiotic concentration. In dotted lines we represent the reversible regime $k_{off},P_{out}\geq k_{on},P_{in}$, in full lines the irreversible regime $k_{off},P_{out}\ll k_{on},P_{in}$. For the irreversible case (full lines), we observe two branches that represent the coexistence of two values of the growth rate, a "large" growth rate and a "near-zero" growth rate. A discontinuity appears when the system jumps from one branch to another. The colors of the curves correspond to different choices of rate constant $k_{B1}$ as shown on the scale on the right. $k_{B,1}$ essentially sets the basal growth rate $\lambda_{0}$ [21] and may vary from one cell to another in a population [35].
(b) Half-inhibition concentration $IC_{50}$ as function of the normalized pre-exposure growth rate in the case of no intermediate steps $m=0$. Symbols represent experimental data points extracted from Ref. [12], which correspond to various antibiotics as shown in the legend.
We recover in Fig.3b the two regimes of antibiotics binding mentioned above, the reversible regime where the half-inhibitory concentration decreases with $\lambda_{0}$ and the irreversible regime where it increases with $\lambda_{0}$. Adding intermediate steps shifts the minimum of the parabola towards lower $\lambda_{0}$ and reduces $IC_{50}$ and thus makes it easier to inhibit growth in the cycle. It also introduces a strong dependence of $IC_{50}$ on the rate constants $k_{1,B}$ in the reversible regime. This reflects that intermediate steps have a stronger impact in reversible pathways as compared to irreversible ones.
Conclusion
In this paper, building on previous works on cellular autocatalytic growth [16], we propose a general model for the inhibition of bacterial growth by antibiotics. This approach goes beyond Ref. [12] because growth laws are no longer introduced as additional constraints and an arbitrary number of steps is introduced in autocatalytic cycles. As we have shown, our model describes well the effects of a large panel of bacteriostatic antibiotics targeting key autocatalytic cycles in E.Coli. We have also found that the two regimes previously identified for ribosome-targeting antibiotics in [12], namely the reversible (strong outflux of inhibitors) and irreversible (small outflux of inhibitors) regimes, should in fact be expected generically for any bacteriostatic inhibitors targeting an autocatalytic cycle.
In the future, we would like to expand our approach towards bacteriocidal antibiotics, which are typically used in conjunction with bacteriostatic antibiotics in a time-dependent manner [36]. To understand cell death, one possibility would be to relate the measure of risk which we have introduced to the extinction probability of the cell. Experiments show significant cell-to-cell heterogeneity in antibiotic susceptibility [37], which require a model for the stochastic growth and death of individual cells and for the fluctuations in population size. In this respect, it is encouraging to see that our model predicts growth bistability, which could cause cell-to-cell heterogeneity, but clearly more work is needed to relate the single-cell and population susceptibility.
Finally, let us also point out that our approach based on autocatalytic cycles is rather general and could be applied beyond cellular biology to other fields, such as ecology [38] or economy, where individuals rather than molecules are able to create more of themselves thanks to autocatalytic cycles but can also be inhibited by toxic agents, either present in their environment or created by themselves as a result of their own growth.
Acknowledgements
We acknowledge inspiring discussions with C. Baroud and E. Maikranz, and insightful comments by L. Dinis and R. Pugatch.